# LEREL: Lipschitz Continuity-Constrained Emotion Recognition Ensemble Learning For Electroencephalography

### v1

@@ 3.2 Experimental Results
In this study, the original data were resampled to 100 Hz to reduce data volume and improve efficiency while retaining key features according to the Nyquist theorem. The resampled data were segmented every 500 samples to extract local features and reduce complexity. The stratified sampling was used to construct training and testing sets, ensuring consistent sample distribution and improving model stability and generalization. The Adam [37] optimizer was employed during model training stage.
On the EAV dataset, we compared LEREL with the reported results in [34], and AMERL [9]. Since AMERL is a multimodal-based emotion recognition model that used active signals (i.e., signals collected when subjects were engaged in conversational stimuli) in its experimental validation, and our method only used EEG signals for learning, we focused on comparing the EEG-based recognition performance of AMERL. LEREL achieved a recognition accuracy of 76.43% and an F1-score of 76.22% on the EAV dataset, significantly outperforming existing methods and demonstrating superior EER performance. The corresponding results are presented in Table 1 and Figure 7.
Table 1: The results of EAV
EAV AMERL-EEG LEREL (ours) EAV AMERL-EEG LEREL (ours)
Subject ACC (%) F1 (%) ACC (%) F1 (%) ACC (%) F1 (%) Subject ACC (%) F1 (%) ACC (%) F1 (%) ACC (%) F1 (%)
1 67.50 - 59.17 - 75.83 75.80 23 63.30 - 50.00 - 76.67 75.67
2 80.00 - 64.17 - 89.17 89.10 24 75.00 - 76.67 - 88.33 88.34
3 52.50 - 54.17 - 80.00 79.97 25 49.20 - 50.00 - 77.50 78.03
4 77.50 - 66.67 - 81.67 81.34 26 58.30 - 49.17 - 73.33 72.92
5 52.50 - 40.00 - 65.00 65.11 27 70.00 - 65.00 - 86.67 86.51
6 47.50 - 48.33 - 79.17 79.35 28 63.30 - 67.50 - 87.50 87.39
7 59.20 - 59.17 - 87.50 87.29 29 51.70 - 42.50 - 62.50 61.96
8 56.70 - 54.17 - 83.33 83.33 30 59.20 - 50.00 - 69.17 68.81
9 44.20 - 45.00 - 78.33 77.81 31 48.30 - 44.17 - 74.17 73.95
10 66.70 - 47.50 - 70.83 70.68 32 60.00 - 55.83 - 71.67 71.36
11 46.70 - 44.17 - 82.50 82.68 33 73.30 - 62.50 - 84.17 84.05
12 65.80 - 45.00 - 79.17 78.13 34 54.20 - 46.67 - 67.50 66.58
13 64.20 - 55.00 - 71.67 72.17 35 30.80 - 28.33 - 44.17 43.93
14 47.50 - $3$ 66.67 65.97 36 58.30 - 60.83 - 79.17 79.37
15 59.20 - 51.67 - 70.83 70.67 37 54.20 - 55.83 - 75.00 75.04
16 55.00 - 42.50 - 71.67 71.81 38 64.20 - 45.83 - 77.50 77.34
17 78.30 - 67.50 - 84.17 84.14 39 59.20 - 50.83 - 80.83 80.46
18 69.20 - 64.17 - 84.17 83.90 40 40.00 - 40.00 - 59.17 58.75
19 55.00 - 51.67 - 75.00 74.95 41 26.70 - 43.33 - 64.17 63.15
20 77.50 - 73.33 - 87.50 87.29 42 72.50 - 65.00 - 84.17 83.89
21 68.30 - 60.00 - 85.83 86.07 Average 59.51 - 53.51 - 76.43 76.22
22 76.70 - 70.83 - 76.67 76.28 - - - - - - -
On the FACED dataset, we evaluated LEREL against the original benchmark results in [35]. LEREL achieved 83.00% recognition accuracy and 72.70% F1-score, substantially exceeding baseline performance metrics and demonstrating enhanced robustness in processing heterogeneous EEG signals. The corresponding results are presented in Table 2 and Figure 7.
Table 2: The results of FACED
Subject ACC(%) Subject ACC(%) Subject ACC(%) Subject ACC(%) Subject ACC(%)
1 92.06 26 84.13 51 72.22 76 82.54 101 86.51
2 87.30 27 76.19 52 86.51 77 88.89 102 85.71
3 81.75 28 85.71 53 93.65 78 84.13 103 80.16
4 88.89 29 98.41 54 81.75 79 89.68 104 79.37
5 66.67 30 88.89 55 95.24 80 87.30 105 75.40
6 73.81 31 86.51 56 95.24 81 87.30 106 91.27
7 87.30 32 80.16 57 90.48 82 80.95 107 78.57
8 79.37 33 64.29 58 80.16 83 93.65 108 84.13
9 89.68 34 72.22 59 95.24 84 78.57 109 84.92
10 85.71 35 93.65 60 90.48 85 83.33 110 83.33
11 72.22 36 86.51 61 65.87 86 76.19 111 77.78
12 65.87 37 73.81 62 90.48 87 90.48 112 86.51
13 84.92 38 85.71 63 82.54 88 84.13 113 85.71
14 80.95 39 90.48 64 92.86 89 89.68 114 78.57
15 76.98 40 71.43 65 95.24 90 76.98 115 84.92
16 72.22 41 86.51 66 69.84 91 67.46 116 96.03
17 71.43 42 66.67 67 77.78 92 84.13 117 87.30
18 79.37 43 84.92 68 81.75 93 80.16 118 81.75
19 74.60 44 80.16 69 90.48 94 81.75 119 81.75
20 89.68 45 69.84 70 87.30 95 86.51 120 82.54
21 86.51 46 87.30 71 93.65 96 91.27 121 95.24
22 85.71 47 81.75 72 82.54 97 89.68 122 80.16
23 80.95 48 81.75 73 80.16 98 100.00 123 75.40
24 48.41 49 95.24 74 87.30 99 95.24 Average 83.00
25 93.65 50 61.11 75 95.24 100 79.37 - -
Figure 5: Confusion Matrix of EAV
Figure 6: Confusion Matrix of FACED
Figure 7: Confusion Matrix of SEED
On the SEED dataset, we further validated the performance of our LEREL. The experimental results showed that LEREL achieved an accuracy of 89.22% and an F1-score of 89.06%, which substantiated the efficacy of LEREL across diverse emotion recognition paradigms. The corresponding results are presented in Table 3 and Figure 7.
Table 3: The results of SEED
Subject ACC(%) F1(%) Subject ACC(%) F1(%) Subject ACC(%) F1(%)
1-1 91.14 92.23 6-1 92.61 89.74 11-1 90.18 87.16
1-2 87.75 85.93 6-2 94.58 94.97 11-2 89.93 89.07
1-3 78.30 80.60 6-3 90.24 92.46 11-3 92.55 92.45
2-1 80.10 78.80 7-1 83.36 86.50 12-1 88.45 91.81
2-2 80.68 83.29 7-2 89.59 89.96 12-2 89.42 88.80
2-3 85.98 86.98 7-3 90.45 88.74 12-3 91.95 91.53
3-1 90.03 89.66 8-1 87.74 88.42 13-1 88.02 86.67
3-2 90.00 87.73 8-2 89.49 91.60 13-2 89.70 91.32
3-3 89.48 85.86 8-3 94.29 91.98 13-3 90.70 91.27
4-1 93.70 88.67 9-1 92.79 94.67 14-1 93.42 93.74
4-2 82.62 80.57 9-2 93.41 90.19 14-2 83.15 83.27
4-3 78.22 79.90 9-3 88.32 85.32 14-3 87.92 85.66
5-1 92.53 92.88 10-1 87.23 87.39 15-1 96.33 97.87
5-2 90.34 92.70 10-2 87.55 85.28 15-2 92.65 94.90
5-3 94.28 94.02 10-3 86.08 87.52 15-3 97.55 97.55
Average 89.22 89.06 - - - - - -
In summary, we conducted rigorous experiments across multiple representative datasets to evaluate LEREL’s performance. These datasets encompass diverse emotional categories, data dimensions, and signal characteristics, establishing a robust evaluation framework. The experimental results show that LEREL significantly outperforms the competing methods in key metrics such as recognition accuracy and F1-score, demonstrating its efficiency, robustness, and wide applicability.
@@ 3.3 Effects of Lipschitz Continuity Constraints
In our study, we applied Lipschitz continuity constraints across key components of the LEREL framework, including the EnhancedNASM component, attention mechanism, and normalization operations, to improve model stability and robustness. Experimental results (Figures 9 and 9) show that these constraints significantly enhanced training stability, particularly on low-accuracy datasets, comfirming their effectiveness for performance improvement.
Figure 8: EAV Sub2 with different Lipschitz constant
Figure 9: EAV Sub35 with different Lipschitz constant
@@ 3.4 Ablation Studies
The proposed LEREL framework employs an ensemble learning strategy comprising four independent sets of feature extractors, enhancers, and classifiers for EEG-based emotion recognition. Each set independently processes EEG signals to improve feature discriminability, while learnable weights optimize classifier integration (Table 4). Experimental results demonstrated that: 1) each classifier achieve relatively high accuracy. 2) the ensemble learning strategy significantly improves stability and robustness, and 3) ablation studies confirm its effectiveness in reducing misclassification rates.
Table 4: The results of ablation studies on the EAV dataset
Ablation model Average accuracy Average F1-score
Classifier 1 69.56 69.24
Classifier 2 68.23 67.90
Classifier 3 61.01 60.29
Classifier 4 62.10 61.72

### v2

@@ 4.3 Experimental Results
In this part, we conducted a comprehensive and systematic evaluation of LEL on the EAV dataset and compared it with the results reported in [46] and those of AMERL (An Multimodal Emotion Recognition Learning) [15]. Unlike our method, which solely relies on EEG signals, AMERL uses active signals collected during conversational stimuli. To ensure a fair comparison, we focused on AMERL’s EEG-based performance.
The experimental results showed that LEL achieved a recognition accuracy of 76.43% and an F1-score of 76.22% on the EAV dataset, surpassing existing methods. When compared to AMERL’s EEG benchmark, LEL demonstrated significant advantages in multiple evaluation metrics, especially in processing EEG data. This indicates that even without multimodal data, LEL’s single EEG signal recognition performance can outperform multimodal approaches. The comparative results are presented in Table 1, and the confusion matrix illustrating LEL’s classification performance on the EAV dataset is shown in Figure 2a.
Figure 2: Confusion Matrices of LEL (Our Model) Across Three Datasets: Left for EAV (a), Center for FACED (b), and Right for SEED (c).
Table 1: Performance Comparison of Different Methods on the EAV Dataset
Dataset Type Method ACC(%) F1(%)
EAV Traditional SVM 48.96 48.67
KNN 50.87 47.68
Decision Tree 45.25 45.28
Others EAV-EEG [46] 59.51 -
AMERL [15] 53.51 -
Ours LEL 76.43 76.22
On the FACED dataset, we conducted a comprehensive evaluation of LEL and compared it with the original benchmark results from [47]. The FACED dataset is renowned for its diversity and complexity, encompassing a wide range of emotional states and EEG signal characteristics, which makes evaluation on this dataset particularly challenging. As shown in Table 2, our LEL demonstrates superior performance on the FACED dataset. Specifically, when faced with various types of EEG signals, LEL exhibited greater adaptability and stability. It could effectively deal with noise and variations in the signals, thereby achieving high-precision emotion recognition. Compared with the original benchmark results from [47], our model achieves a recognition accuracy of 83.00% and an F1-score of 72.70% . When contrasted with other methods like ACCNet [49] which achieved 63.07% ACC and 67.99% F1-score, FC_STGNN [50] with 61.12% ACC and 61.84% F1-score, EEGNet [51] which obtained 61.55% ACC and 60.61% F1-score, LGGNet [52] with 54.43% ACC and 59.26% F1-score, BF-GCN [53] which got 55.23% ACC and 63.67% F1-score, and FBCNet [54] with 63.14% ACC and 62.93% F1-score, our LEL significantly outperforms them. The detailed comparative results, including various performance metrics, are presented in Table 2. Additionally, Figure 2b visually illustrates LEL’s classification performance on the FACED dataset in the form of a confusion matrix.
Table 2: Performance Comparison of Different Methods on the FACED Dataset
Dataset Type Method ACC(%) F1(%)
FACED Traditional SVM 49.05 47.38
KNN 45.44 44.63
Decision Tree 41.46 41.71
Others ACCNet [49] 63.07 67.99
FC_STGNN [50] 61.12 61.84
EEGNet [51] 61.55 60.61
LGGNet [52] 54.43 59.26
BF-GCN [53] 55.23 63.67
FBCNet [54] 63.14 62.93
Ours LEL 83.00 72.70
On the SEED dataset, we further comprehensively evaluated LEL. The SEED dataset, a key benchmark for emotion recognition, presents a challenging environment for assessing model performance. When compared with other methods such as ACCNet [49] which achieved 80.50% ACC and 79.43% F1-score, FC_STGNN [50] with 69.50% ACC and 68.22% F1-score, EEGNet [51] which obtained 74.50% ACC and 73.93% F1-score, LGGNet [52] with 72.24% ACC and 71.96% F1-score, BF-GCN [53] which got 68.83% ACC and 67.91% F1-score, and FBCNet [54] with 77.31% ACC and 78.26% F1-score, our LEL significantly outperforms them by achieving an accuracy of 87.22% and an F1-score of 87.06%. LEL showed stable recognition precision across all emotion categories, without significant performance fluctuations, indicating its excellent stability and adaptability to data changes under different emotion states.The detailed comparative results are presented in Table 3, and Figure 2c intuitively shows the confusion matrix of LEL’s classification performance.
Table 3: Performance Comparison of Different Methods on the SEED Dataset
Dataset Type Method ACC(%) F1(%)
SEED Traditional SVM 66.84 66.73
KNN 63.50 67.22
Decision Tree 66.41 66.38
Others ACCNet [49] 80.50 79.43
FC_STGNN [50] 69.50 68.22
EEGNet [51] 74.50 73.93
LGGNet [52] 72.24 71.96
BF-GCN [53] 68.83 67.91
FBCNet [54] 77.31 78.26
Ours LEL 87.22 87.06
Experimental results demonstrate LEL’s robust performance across benchmark datasets: on EAV, it achieved 76.43% accuracy and 76.22% F1-score; on FACED, 83.00% accuracy; on SEED, 87.22% accuracy and 87.06% F1-score. These metrics confirm the framework’s efficacy in capturing complex emotional features and its reliability for diverse emotion recognition tasks.
@@ 4.4 Effects of Lipschitz Continuity Constraints
4.4.1 Analysis of Training Process
In this study, we introduce Lipschitz continuity constraints into key components of the LEL framework. Lipschitz continuity is a mathematical property that bounds the rate of change of a function, ensuring that small variations in input do not lead to disproportionately large changes in output. In the context of deep learning, enforcing Lipschitz constraints serves to regulate model complexity, mitigate overfitting, and enhance robustness against noise and perturbations—factors that are particularly critical in emotion recognition tasks, where data is inherently noisy, variable, and susceptible to outliers.
By embedding Lipschitz constraints within the LEL architecture, we effectively controlled the magnitude of weight updates during training. This led to reduced fluctuations and facilitated smoother convergence. As illustrated in Figures 4 and 4, our experimental results demonstrate that these constraints significantly improve training stability, especially on low-accuracy datasets characterized by noisy signals and complex feature distributions. Such stability is essential not only for optimizing LEL during training but also for ensuring its reliability in real-world applications.
Figure 3: EAV Sub2 with different Lipschitz constant.
Figure 4: EAV Sub35 with different Lipschitz constant.
4.4.2 Analysis with ROC Curve
The ROC curve is a widely used tool for evaluating the performance of classification models. It illustrates the trade-off between the True Positive Rate (TPR) and False Positive Rate (FPR) across various decision thresholds, thereby providing a clear visualization of a model’s discriminative capability. A larger Area Under the Curve (AUC) value indicates stronger class separation and better overall classification performance. Experimental results reveal that, as the number of training epochs increases, the ROC curves for different emotional categories progressively shift toward the upper-left corner, accompanied by a rising trend in AUC values. This suggests that extended training allows the model to better capture complex patterns and subtle distinctions in the data, thereby enhancing both classification accuracy and reliability.
In this context, the Lipschitz constant plays a critical role in regulating model complexity, preventing overfitting, and controlling the degree of adaptation to the training data. For models trained with fewer epochs, selecting an appropriate Lipschitz constant is particularly important. A smaller Lipschitz constant constrains model complexity, enabling the acquisition of stable feature representations during the early training phase (Figure 5a). This facilitates the learning of fundamental data patterns, preserving a baseline level of emotional feature discrimination. A moderate Lipschitz constant strikes a balance between overfitting prevention and sufficient learning capacity (Figure 5b). It allows the model to more effectively capture emotional features with a moderate number of training epochs, resulting in steeper ROC curves and higher AUC values. Conversely, a larger Lipschitz constant permits the model to learn more intricate feature representations (Figure 5c), which can further steepen the ROC curve and increase the AUC. However, this also heightens the model’s sensitivity to noise, potentially increasing the risk of overfitting—especially when the number of training epochs is high.
Figure 5: The figure presents nine line charts visualizing the ROC curves for emotion classification under different training parameters: (a)-(c) 30 training epochs with Lipschitz constants 0.1, 1.0, and 10.0 (d)-(f) 50 training epochs with Lipschitz constants 0.1, 1.0, and 10.0 (g)-(i) 100 training epochs with Lipschitz constants 0.1, 1.0, and 10.0 Each chart displays ROC curves for different emotion categories including Joy, Tenderness, Inspiration, Amusement, Anger, Disgust, Fear, Sadness, and Neutral. The curves illustrate the trade-off between true positive rate and false positive rate for each emotion classifier.
From the perspectives of epochs and Lipschitz constant, they jointly influence the training process and final performance of the model. When training starts and epochs are few, choosing a smaller Lipschitz constant helps the model quickly and stably learn basic data features, avoiding overfitting and laying a solid foundation for subsequent training. As training progresses and epochs increase, the Lipschitz constant can be appropriately increased to enable the model to learn more complex feature representations and further enhance classification performance. However, when adjusting the Lipschitz constant, it is crucial to monitor the model’s generalization ability to prevent overfitting due to excessive model complexity. Reasonable epoch planning and Lipschitz constant selection allow the model to continuously optimize during training with appropriate learning speed and feature extraction ability. This ultimately leads to better classification results on test data. In emotion classification tasks, which are complex and prone to noise interference, this balance and optimization are especially important.
From the figures, it can be observed that the ROC curves under different epochs and Lipschitz constants show the changing trend of model performance. When both epochs and Lipschitz constant are small, the curve is relatively flat, indicating limited learning ability and weak emotional feature distinction at this stage (Figure 5a). As epochs increase, the curve gradually moves toward the upper left corner, and model classification performance improves significantly (Figure 5d). At the same time, appropriately increasing the Lipschitz constant can make the curve steeper and further enhance classification performance (Figure 5e and 5f). However, when the Lipschitz constant is too large, the curve may fluctuate (Figure 5h and 5i), indicating the model is starting to overfit the training data, leading to a decline in generalization ability. Therefore, in practical applications, it is necessary to comprehensively consider the settings of epochs and Lipschitz constant. Through experiments and validation, the optimal combination should be found to achieve the best model performance.
4.4.3 Visualization Using t-SNE
To analyze and compare how different training iterations and Lipschitz constants affect the model, we conducted experiments on the FACED dataset. This dataset contains nine 9 emotion classes, offering greater complexity and diversity that better demonstrates the experimental results. The dataset covers a wide range of emotions, including joy, gentleness, inspiration, entertainment, anger, disgust, fear, sadness, and neutrality. These are representative of human daily emotions and offer comprehensive data for emotion analysis studies. We changed the training epochs (30, 50, and 100) to see how training times impact model performance. We also used Lipschitz constants (0.1, 1.0, and 10.0) as control variables to study their influence on the model’s generalization and feature learning. Lipschitz constants limit model complexity and affect how well the model fits the training data, helping prevent overfitting and improve robustness.
Figure 6f shows the t-SNE visualization of raw EEG signals. The different emotion categories show some clustering in 2D space, but the boundaries between them are unclear. This means the emotion features in raw EEG signals are complex and overlapping, making accurate classification from raw signals challenging. This stresses the need for effective feature extraction and model training to better separate emotion features and improve classification accuracy. t-SNE visualization helps us see the distribution of emotion features in raw EEG signals, providing a basis for further model training and optimization.
Figure 6a, 6b and 6c show the t-SNE visualization after 30 epochs. The model’s output indicates improved emotion category clustering, with some emotions (like joy and gentleness) forming tighter clusters. This suggests the model has initial learning capacity for these features. However, the boundaries are still unclear with many mixed data points, indicating insufficient feature learning due to limited training. As training increases to 50 epochs (Figure 6d), clustering improves and boundaries become clearer. For example, fear and sadness form more distinct clusters, showing better feature differentiation. This means the model has learned more about EEG emotion features, with more accurate representations and improved separation. At 100 epochs (Figure 6e) , the model shows nearly ideal emotion category distribution. Boundaries are clear, and most categories form tight, independent clusters, especially neutral versus negative emotions. This proves sufficient training enhances model performance, enabling accurate EEG emotion feature extraction and classification.
Figure 6: The figure presents six scatter plots visualizing emotion classification results under different training parameters: (a) 30 training epochs with Lipschitz constant 0.1 (b) 30 training epochs with Lipschitz constant 1.0 (c) 30 training epochs with Lipschitz constant 10.0 (d) 50 training epochs with Lipschitz constant 0.1 (e) 100 training epochs with Lipschitz constant 0.1 (f) Visualization of raw EEG signals Each plot shows data points color-coded by emotion categories including Joy, Tenderness, Inspiration, Amusement, Anger, Disgust, Fear, Sadness, and Neutral. The visualizations demonstrate how different training configurations affect the separation and clustering of emotional states in the feature space.
Figure 6a shows the t-SNE visualization with a Lipschitz constant of 0.1. The model’s output shows limited emotion category clustering. The small constant restricts model complexity, leading to stable feature representations even with fewer training epochs and clearer emotion boundaries. This helps avoid overfitting and preserves differentiation ability, which is crucial for emotion classification due to the subtlety and noise sensitivity of EEG emotion features.With the Lipschitz constant set to 1.0 (Figure 6b), the model balances complexity and stability. Under moderate training epochs, the clustering and boundary clarity of emotion categories are significantly improved, demonstrating excellent classification performance. This indicates that a moderate Lipschitz constant can effectively control the model’s learning behavior. It enables the model to thoroughly learn the emotional features in EEG signals without overfitting, thereby enhancing classification accuracy. During this phase, the model can better capture the subtle details of emotional features while maintaining effective differentiation between various emotion categories.When the Lipschitz constant is increased to 10.0 (Figure 6c), the model is permitted to learn more complex feature representations. With fewer training epochs, the clustering of certain emotion categories may not be tight and could display some dispersion. However, as the number of training iterations increases, the model can capture more nuanced emotional features, further enhancing the classification outcomes.
This suggests that a larger Lipschitz constant can assist the model in learning richer features and improving classification performance when there is adequate training. Nevertheless, a higher Lipschitz constant might also make the model more sensitive to noise in the training data, thereby increasing the risk of overfitting. Consequently, when employing a larger Lipschitz constant, it is essential to carefully balance the model’s complexity with its generalization capabilities. Overall, setting the Lipschitz constant appropriately is crucial for the model’s emotional classification performance. With fewer training epochs, a smaller Lipschitz constant aids the model in achieving stable feature representations and yielding favorable classification results. On the other hand, under sufficient training conditions, moderately increasing the Lipschitz constant allows the model to learn more diverse features, thereby further elevating its performance.
@@ 4.5 Ablation Studies
The LEL proposed in this study employs an integrated learning strategy with four sets of feature extractors, enhancers, and classifiers to process EEG signals independently and enhance feature distinctiveness. A learnable weight mechanism combines these classifiers to boost overall performance. As shown in Table 4, each individual classifier achieves a relatively high accuracy: Classifier 1 at 69.56%, Classifier 2 at 68.23%, Classifier 3 at 61.01%, and Classifier 4 at 62.10%. When these classifiers are integrated, the average accuracy of the full model rises significantly to 74.25%, with the best - case accuracy reaching 76.43%, and the F1 - score improving to 73.94% and 76.22% respectively. This highlights the integrated strategy’s effectiveness in boosting stability, robustness, and reducing misclassification.
Further analysis reveals that the integrated learning strategy not only mitigates overfitting by aggregating predictions from multiple classifiers but also enhances the model’s adaptability to noisy and complex emotional features. This enables the LEL to deliver more reliable outputs in complex emotional recognition tasks. The learnable weight mechanism for classifier integration allows the model to automatically adjust each classifier’s weight in the final decision - making process, further enhancing overall performance and ensuring good generalization across different datasets and emotional categories.
In summary, the integrated learning strategy in the LEL significantly enhances the accuracy and reliability of emotion recognition. The excellent performance of each individual classifier, combined with the optimization from the integrated strategy, offers a novel approach for EEG-based emotion recognition with a promising future outlook. Future work will explore applying this method to more emotion recognition tasks to evaluate its potential and value in diverse scenarios.
Table 4: The results of ablation studies on the EAV positive dataset.
Ablation model Average ACC(%) Average F1(%)
Global Feature Extractor 69.56 69.24
Channel Energy Extractor 68.23 67.90
Channel Feature Mixer 61.01 60.29
Band Magnitude Mixer 62.10 61.72